# Efficient Diffusion Models: A Survey

### v1

@@ 3.3.3 Truncated Sampling
Truncated sampling enhances the efficiency of sample generation in diffusion models by strategically reducing redundant computations, thereby lowering computational costs while maintaining high-quality outputs. Methods like Early Exit focus on skipping unnecessary computations in later stages of the diffusion process when predictions are confident. Meanwhile, Retrieval-Guided Initialization improves efficiency in the early stages by leveraging retrieved examples to provide a better initialization, effectively skipping parts of the iterative refinement process. These approaches allocate computation more effectively by focusing resources on the most critical steps of the sampling process.
Early Exit. Early exit is a technique to improve the efficiency of large neural networks by reducing unnecessary computation. It allows intermediate layers to make early predictions, saving computational resources while maintaining performance. The core idea is to use intermediate predictions when they are confident enough, thereby skipping further layers for simpler inputs. Schuster et al. (2021) introduce Confident Adaptive Transformers (CATs). As shown in Figure 18, model $\mathcal{G}(X)$ provably consistent with the original $\mathcal{F}$ with arbitrarily high probability. They enhance transformer efficiency by adding prediction heads at intermediate layers and using a meta classifier to decide when to stop computation. This ensures high confidence in predictions, significantly improving efficiency on NLP tasks without sacrificing accuracy. Formally, they create a model $\mathcal{G}(X)$ that includes early classifiers $\mathcal{G}=\{\mathcal{F}_{1},\mathcal{F}_{2},\dots,\mathcal{F}_{l}\},\quad\text{ensuring}\quad P(\mathcal{G}(x)=\mathcal{F}(x))\geq 1-\epsilon$.
Figure 18: The CAT model $\mathcal{G}$ Schuster et al. (2021) improves computational efficiency by enabling early exits on certain inputs while ensuring predictive consistency with the full model $\mathcal{G}$.
Building on this, another work Moon et al. (2024) propose Adaptive Score Estimation (ASE) for diffusion models. ASE reduces sampling time by optimizing computation allocation at different time steps, introducing a time-dependent exit schedule. This method improves sampling efficiency while preserving image quality. The key idea is to dynamically allocate computation based on the difficulty of score estimation at each time step, allowing for early exits in simpler cases.
Retrieval-Guided Initialization. Retrieval-Guided Initialization combines the efficiency of retrieval mechanisms with the generative power of diffusion models. As illustrated in Figure 19, this approach begins with a retriever that selects relevant images from a database based on the input text prompts. These retrieved images, organized by specific keywords, serve as contextual guidance for the diffusion model, enhancing the relevance and quality of the generated output image.Blattmann et al. (2022) introduce a semi-parametric model that integrates neural network-based generative models with a retrieval mechanism, guided by examples from a large dataset. This approach shows the feasibility of combining retrieval with neural synthesis, although it requires a large dataset.  Sheynin et al. (2022) propose integrating a K-Nearest Neighbors (KNN) retrieval mechanism with diffusion models. By retrieving similar images from a large dataset, the model effectively guides the diffusion process, helping to generate high-quality images efficiently. This approach addresses the challenge of generating high-quality images with limited resources. Chen et al. (2022b) introduce a retrieval-augmented framework specifically for text-to-image generation, retrieving relevant images based on textual descriptions to guide the diffusion process and align the output with the input text.
Finally, Zhang et al. (2023b) introduces a learning-free inference mechanism for diffusion models. Instead of training complex networks, they retrieve and reuse diffusion trajectories from a precomputed database, significantly reducing the computational burden and providing an efficient solution for real-time applications.
@@
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=left, text centered, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=18em,font=,, where level=2text width=15em,font=,, where level=3text width=24em,font=,, [ Compression, ver [ Quantization, fill=blue!10 [ Post-Training Quantization, fill=blue!10 [ PTQ4DM (Shang et al., 2023), Q-Diffsuion (Li et al., 2023), BRECQ (Li et al., 2021),
RAQ (Kim et al., 2024b), PTQD (He et al., 2024), leaf, text width=45em ] ] [ Quantization-Aware Training, fill=blue!10 [ TDQ (So et al., 2024), QALoRA (He et al., 2023), leaf, text width=45em ] ] ] [ Pruning, fill=blue!10 [ Diff-Pruning (Fang et al., 2023), LD-Pruner (Castells et al., 2024), LayerMerge (Kim et al., 2024a), LAPTOPDiff (Zhang et al., 2024), leaf, text width=61.7em ] ] [ Knowledge Distillation, fill=blue!10 [ Vector Field Distillation, fill=blue!10 [ Denoising Student (Luhman & Luhman, 2021), Progressive Distillation (Salimans & Ho, 2022),
 Meng et al. (2023), CM (Song et al., 2023), leaf, text width=45em ] ] [ Generator Distillation, fill=blue!10 [ SDS (Poole et al., 2022), VSD (Wang et al., 2024b), Diff-Instruct (Luo et al., 2024),
CSD (Decatur et al., 2024), leaf, text width=45em ] ] ] ]
Figure 20: Summary of compression techniques for DMs.
3.4 Compression
Model compression enhances efficiency by reducing the sizes and the amount of arithmetic operations of DM. As summarized in Figure 28, model compression techniques for DMs can be grouped into three categories: quantization, pruning, and knowledge distillation. These three categories are orthogonal to each other, and compress DMs from different perspectives.
3.4.1 Quantization
Figure 21: Illustrations of the quantization.
Quantization-Aware Training. Different from PTQ, QAT quantizes diffusion models during the training process, allowing models to learn quantization-friendly representations. Since QAT requires additional training after introducing quantization operators, it is much more expensive and time-consuming than PTQ. So et al. (2024) proposes a novel quantization method that enhances output quality by dynamically adjusting the quantization interval based on time step information. The proposed approach integrates with the Learned Step Size Quantization (Esser et al., 2019) framework, replacing the static quantization interval with a dynamically generated output from the Time-Dynamic Quantization module. This dynamic adjustment leads to significant improvements in the quality of the quantized outputs. He et al. (2023) introduces a quantization-aware low-rank adapter that integrates with model weights and is jointly quantized to a low bit-width. This approach distills the denoising capabilities of full-precision models into their quantized versions, utilizing only a few trainable quantization scales per layer and eliminating the need for training data.
3.4.2 Pruning
Pruning compresses DMs by removing redundant or less important model weights. Currently, most pruning methods for DMs focus on pruning structured patterns such as groups of consecutive parameters or hierarchical structures. For instance, Diff-Pruning (Fang et al., 2023) introduces the first dedicated method designed for pruning diffusion models. Diff-Pruning leverages Taylor expansion over pruned timesteps to estimate the importance of weights. By filtering out non-contributory diffusion steps and aggregating informative gradients, Diff-Pruning enhances model efficiency while preserving essential features.
LD-Pruner (Castells et al., 2024), as illustrated in Figure 23, on the other hand, proposes a pruning method specifically designed for Latent Diffusion Models (LDMs) The key innovation of LD-Pruner lies in its utilization of the latent space to guide the pruning process. The method enables a precise assessment of pruning impacts by generating multiple sets of latent vectors—one set for the original Unet and additional sets for each modified Unet where a single operator is altered. The importance of each operator is then quantified using a specialized formula that considers shifts in both the central tendency and variability of the latent vectors. This approach ensures that the pruning process preserves model performance while adapting to the specific characteristics of LDMs.
Figure 24: Illustrations of the pruning.
Lastly, LAPTOPDiff (Zhang et al., 2024) introduces a layer-pruning technique aimed at automatically compressing the U-Net architecture of diffusion models. The core of this approach is an effective one-shot pruning criterion, distinguished by its favorable additivity property. This property ensures that the one-shot performance of the pruning is superior to other traditional layer pruning methods and manual layer removal techniques. By framing the pruning problem within the context of combinatorial optimization, LAPTOPDiff simplifies the pruning process while achieving significant performance gains. The proposed method stands out for its ability to provide a robust one-shot pruning solution, offering a clear advantage in compressing diffusion models efficiently.
@@ 3.4.3 Knowledge Distillation
Vector Field Distillation. Vector field distillation improves the efficiency of deterministic sampling in diffusion models by transforming the generative ODE into a new generative vector field. This approach reduces the number of NFEs needed to produce samples of similar quality. Luhman & Luhman (2021) first proposes a strategy to distill a DDIM sampler into a Gaussian model that needs only one NFE for sampling. In this approach, a conditional Gaussian model serves as the student model, and the training process involves minimizing the conditional KL divergence between this student model and the DDIM sampler.
@@
While this method advances the application of knowledge distillation to diffusion models, it still has computational inefficiencies, as it necessitates generating the final outputs of DDIM or other ODE samplers, which entails hundreds of NFEs for each training batch. Salimans & Ho (2022) proposes a progressive distillation strategy to train a student model to use half the NFEs of the teacher model by learning its two-step prediction strategy, as illustrated in Figure 26. Once the student model accurately predicts the teacher’s two-step sampling strategy, it replaces the teacher model, and a new student model is trained to further reduce the sampling steps. This method reduces the NFEs significantly, achieving 250 times greater efficiency with only a 5% drop in performance.
A two-stage distillation strategy is proposed by Meng et al. (2023) to address the challenge of transferring knowledge from classifier-free guided conditional diffusion models like DALL·E-2 (Ramesh et al., 2022) and Stable Diffusion (Rombach et al., 2022). In the first stage, a student model is trained with classifier-free guidance to learn from the teacher diffusion model. The second stage employs the progressive diffusion strategy to further reduce the number of diffusion steps for the student model. This two-stage approach is applied to both pixel-space and latent-space models for various tasks, including text-guided generation and image inpainting.  Song et al. (2023) introduce the Consistency Model (CM), which leverages the self-consistency property of generative ODEs in diffusion models. Instead of directly mimicking the output of the generative ODE, their method focuses on minimizing the difference in the self-consistency function. By randomly diffusing a real data sample and simulating a few steps of the generative ODE to generate another noisy sample on the same ODE path, the model inputs these two noisy samples into a student model.
@@ 4 System-Level Efficiency Optimization
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=left, text centered, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=22em,font=,, where level=2text width=15em,font=,, where level=3text width=24em,font=,, [ System, ver [ Optimized Hardware-Software Co-design, fill=blue!10 [  Chen et al. (2023c), SDA (Yang et al., 2023a),  Choi et al. (2024a), leaf, text width=61.7em ] ] [ Parallel Computing, fill=blue!10 [ DistriFusion (Li et al., 2024a), PipeFusion (Wang et al., 2024a), SwiftDiffusion (Li et al., 2024b), DiffusionPipe (Tian et al., 2024), leaf, text width=61.7em ] ] [ Caching Technique, fill=blue!10 [ NIRVANA (Agarwal et al., 2024), DeepCache (Ma et al., 2024b), Block Caching (Wimbauer et al., 2024), L2C (Ma et al., 2024a),
FORA (Selvaraju et al., 2024), leaf, text width=61.7em ] ] ]

### v2

@@ 3.3.3 Truncated Sampling
Truncated sampling enhances the efficiency of sample generation in diffusion models by strategically reducing redundant computations, thereby lowering computational costs. This optimization category encompasses several approaches, with Early Exit and Retrieval-Guided Initialization representing two primary strategies that target different phases of the diffusion process to improve computational efficiency. Specifically, early Exit focuses on terminating unnecessary computations in later stages of the diffusion process when predictions are confident. Conversely, Retrieval-Guided Initialization improves efficiency in the early stages by leveraging retrieved examples to provide a better initialization, effectively bypassing parts of the iterative refinement process. These approaches allocate computation more effectively by focusing resources on the most critical steps of the sampling process.
Early Exit. Recent papers that focus on early exit mechanisms have gained significant attention in the Large Language Models (LLMs) domain (Schuster et al., 2021). By implementing early exit strategies, these methods bypass unnecessary computations in deeper layers when outputs can be generated at earlier stages, thereby substantially reducing inference time and computational resources.
Similarly, for computation-intensive diffusion models that typically require numerous denoising steps, early exit techniques present a promising approach to accelerate the generation process dramatically.  Moon et al. (2024a) proposes a simple yet effective early exiting framework called Adaptive Score Estimation (ASE) specifically designed to accelerate the sampling process of diffusion models. The method is based on the key observation that score estimation difficulty varies across different timesteps in the diffusion process, with timesteps closer to the noise distribution requiring fewer computational resources than those closer to the data distribution.
The schematic in Figure 18 illustrates the time-dependent exit schedule used in ASE. As shown in the figure, the model progressively skips more building blocks as the diffusion process moves closer to the noise distribution, where the score estimation becomes easier. For example, in the DiT model (Peebles & Xie, 2023), the blocks are dropped progressively as the timestep approaches 1, closer to the noise regime. Conversely, when the timestep is closer to the data regime, more blocks are retained to ensure accurate score estimation. Moreover, the U-ViT model (Bao et al., 2023) follows a similar dropping schedule, but due to the long skip connections between the encoder and decoder, only decoder blocks are skipped. This preserves critical information from the encoder while still speeding up the process.
Figure 18: Illustration of the time-dependent exit schedule, where more blocks are skipped as the sampling process moves toward the noise distribution, optimizing computational efficiency Moon et al. (2024a).
Through carefully designed time-varying exit schedules, the method significantly accelerates sampling speed while preserving generation quality. It can also be seamlessly integrated with various diffusion model solvers to enhance overall efficiency further.
In contrast,  Tang et al. (2023a) introduces DeeDiff, which approaches early exiting through a dynamic uncertainty-aware framework. While ASE relies on static, predefined exit schedules that are fixed during inference, DeeDiff incorporates a timestep-aware uncertainty estimation module (UEM) that adaptively assesses the prediction uncertainty of each intermediate layer at every sampling step. The UEM directly estimates uncertainty values from the features at each layer using lightweight, fully-connected networks, which are trained to indicate how well the current layer’s output matches the final layer’s prediction. During inference, when a layer’s uncertainty falls below a predefined threshold, the model exits early at that layer, bypassing deeper layers for that particular timestep. Despite their different implementation strategies, both methods demonstrate that early exiting frameworks can effectively balance computational efficiency and generation quality, offering practical solutions for deploying large-scale diffusion models.
Retrieval-Guided Denosing. Retrieval-Guided Initialization combines the efficiency of retrieval mechanisms with the generative power of diffusion models, and can be applied across various generative tasks. This method guides the generation process by retrieving samples or data relevant to the input. For example, as illustrated in Figure 19, when the input is a text prompt, the retriever selects relevant images from a database, which then serve as contextual guidance for the diffusion model to generate a coherent output. By providing a more informed state for the diffusion model, this approach helps the model converge more quickly and generate higher-quality outputs.
Retrieval-Guided Denoising is widely used in text-to-image diffusion tasks, offering an efficient way to generate images that align with textual descriptions. The kNN-Diffusion (Sheynin et al., 2022) method integrates large-scale retrieval techniques with the generative power of diffusion models, offering an efficient approach for image generation. The key idea is to use the CLIP (Radford et al., 2021) image encoder to map images into a shared embedding space and employ the k-Nearest-Neighbors (kNN) retrieval to identify the k most similar image embeddings. Notably, this method does not require paired text-image datasets, which is a significant advantage for domains where such datasets are scarce. During training, the model is conditioned on image embeddings, and kNN retrieval is used to extend the distribution of conditioning embeddings, which helps the model generalize better and bridge the gap between text and image distributions. In the sampling phase, the model receives text input, which is converted to text embedding using CLIP. The kNN retrieval mechanism is then applied to find the k most similar image embeddings based on the text embedding. These retrieved image embeddings serve as additional conditional information, guiding the generation process and ensuring that the generated image aligns with the textual description. Consequently, kNN-Diffusion improves both the efficiency of image generation and the computational resource consumption, making it a highly efficient solution for text-to-image generation tasks. Unlike KNN-Diffusion, which primarily relies on image retrieval to guide the sampling process, Re-Imagen (Chen et al., 2022b) enhances the generation process by incorporating multimodal retrieval to balance computational cost and output quality, using both image and text pairs. Re-Imagen’s retrieval method is based on an external multimodal knowledge base. During sampling, given an input text, the model queries an external database to retrieve the most relevant image-text pairs. These retrieved pairs are then used as additional conditional inputs for the generation process. To identify the most relevant pairs, the model employs either BM25 (Robertson et al., 2009) or CLIP (Radford et al., 2021) similarity measures, which evaluate the alignment between the input text and the image-text pairs in the database. Therefore, this approach provides more diverse visual information, especially for rare or unseen entities. Additionally, Re-Imagen employs a cascaded diffusion architecture (Ho et al., 2022a), which allows it to generate high-quality images more efficiently compared to kNN-Diffusion, as it reduces the computational cost by progressively refining images at different resolutions.
@@
Besides approaches that rely on similarity measures like CLIP embeddings to retrieve images or texts, ReDi (Retrieval-based Diffusion) (Zhang et al., 2023c) offers a fundamentally different approach to accelerating diffusion model inference. ReDi directly retrieves precomputed trajectories from the diffusion process itself. The method constructs a knowledge base of sample trajectories where each entry contains a key-value pair. For instance, in the forward diffusion process (noise-adding process), an early state sample $x_{k}$ serves as the key, while a later state sample $x_{v}$ serves as the value. During inference, ReDi first generates the initial steps of a trajectory up to $x_{k}^{\prime}$, uses this as a query to find a similar state $x_{k}$ in the knowledge base, and then skips intermediate computational steps by jumping to the retrieved $x_{v}$ before continuing the generation process. By skipping a large portion of intermediate steps, ReDi significantly reduces the number of function estimations (NFEs) required during sampling. Experiments with Stable Diffusion demonstrate that ReDi achieves a two times speedup with comparable quality and enables effective zero-shot domain adaptation for tasks like image stylization without requiring domain-specific knowledge bases.
@@ 3.3.4 Knowledge Distillation
Vector Field Distillation. Vector field distillation improves the efficiency of deterministic sampling in diffusion models by transforming the generative ODE into a new generative vector field. This approach reduces the number of NFEs needed to produce samples of similar quality. Luhman & Luhman (2021) first proposes a strategy to distill a DDIM sampler into a Gaussian model that needs only one NFE for sampling. In this approach, a conditional Gaussian model serves as the student model, and the training process involves minimizing the conditional KL divergence between this student model and the DDIM sampler. While this method advances the application of knowledge distillation to diffusion models, it still has computational inefficiencies, as it necessitates generating the final outputs of DDIM or other ODE samplers, which entails hundreds of NFEs for each training batch. Salimans & Ho (2022) proposes a progressive distillation strategy to train a student model to use half the NFEs of the teacher model by learning its two-step prediction strategy, as illustrated in Figure 21. Once the student model accurately predicts the teacher’s two-step sampling strategy, it replaces the teacher model, and a new student model is trained to further reduce the sampling steps. This method reduces the NFEs significantly, achieving 250 times greater efficiency with only a 5% drop in performance.
@@
A two-stage distillation strategy is proposed by Meng et al. (2023) to address the challenge of transferring knowledge from classifier-free guided conditional diffusion models like DALL·E-2 (Ramesh et al., 2022) and Stable Diffusion (Rombach et al., 2022). In the first stage, a student model is trained with classifier-free guidance to learn from the teacher diffusion model. The second stage employs the progressive diffusion strategy to further reduce the number of diffusion steps for the student model. This two-stage approach is applied to both pixel-space and latent-space models for various tasks, including text-guided generation and image inpainting.
Song et al. (2023) firtstly introduces the Consistency Model (CM), which leverages the self-consistency property of generative ODEs in diffusion models. Instead of directly mimicking the output of the generative ODE, their method focuses on minimizing the difference in the self-consistency function. By randomly diffusing a real data sample and simulating a few steps of the generative ODE to generate another noisy sample on the same ODE path, the model inputs these two noisy samples into a student model. Consequently, Chen et al. (2025) accelerates consistency models via hybrid distillation, enabling 1-4 step generation. It transforms pre-trained flow models into TrigFlow (Lu & Song, 2024) without retraining, preserving trajectory alignment while boosting fidelity via adversarial training LADD.
Wu et al. (2024) proposes Multi-scale Latent Point Consistency Model (MLPCM), which implements one-step generation through consistency distillation, combines multi-scale latent space and 3D attention mechanisms to reduce computational complexity. Moreover,  Yin et al. (2024) proposes Distribution Matching Distillation (DMD), which distills multi-step diffusion models into a single-step generator by introducing a distribution-level matching objective that minimizes the KL divergence between real and synthetic data distributions, alongside a regression loss to align large-scale structural features.
@@
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=left, text centered, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=7em,font=,, where level=2text width=15em,font=,, where level=3text width=24em,font=,, [ Compression, ver [ Quantization, fill=blue!10 [ Post-Training Quantization, fill=blue!10 [ PTQ4DM (Shang et al., 2023), Q-Diffsuion (Li et al., 2023), BRECQ (Li et al., 2021),
RAQ (Kim et al., 2024b), PTQD (He et al., 2024), leaf, text width=43em ] ] [ Quantization-Aware Training, fill=blue!10 [ TDQ (So et al., 2024), QALoRA (He et al., 2023), leaf, text width=43em ] ] ] [ Pruning, fill=blue!10 [ Diff-Pruning (Fang et al., 2023), LD-Pruner (Castells et al., 2024), LayerMerge (Kim et al., 2024a), LAPTOPDiff (Zhang et al., 2024a), leaf, text width=59.7em ] ] ]
Figure 23: Summary of compression techniques for DMs.
3.4 Compression
Model compression enhances efficiency by reducing the sizes and the amount of arithmetic operations of DM. As summarized in Figure 28, model compression techniques for DMs can be grouped into quantization and pruning. These two categories are orthogonal to each other, and compress DMs from different perspectives.
Figure 24: Illustrations of the quantization.
3.4.1 Quantization
Quantization-Aware Training. Different from PTQ, QAT quantizes diffusion models during the training process, allowing models to learn quantization-friendly representations. Since QAT requires additional training after introducing quantization operators, it is much more expensive and time-consuming than PTQ.
So et al. (2024) proposes a novel quantization method that enhances output quality by dynamically adjusting the quantization interval based on time step information.
Figure 27: Illustrations of the pruning.
The proposed approach integrates with the Learned Step Size Quantization (Esser et al., 2019) framework, replacing the static quantization interval with a dynamically generated output from the Time-Dynamic Quantization module. This dynamic adjustment leads to significant improvements in the quality of the quantized outputs. He et al. (2023) introduces a quantization-aware low-rank adapter that integrates with model weights and is jointly quantized to a low bit-width. This approach distills the denoising capabilities of full-precision models into their quantized versions, utilizing only a few trainable quantization scales per layer and eliminating the need for training data.
3.4.2 Pruning
Pruning compresses DMs by removing redundant or less important model weights. Currently, most pruning methods for DMs focus on pruning structured patterns such as groups of consecutive parameters or hierarchical structures. For instance, Diff-Pruning (Fang et al., 2023) introduces the first dedicated method designed for pruning diffusion models. Diff-Pruning leverages Taylor expansion over pruned timesteps to estimate the importance of weights. By filtering out non-contributory diffusion steps and aggregating informative gradients, Diff-Pruning enhances model efficiency while preserving essential features. LD-Pruner (Castells et al., 2024), as illustrated in Figure 26, on the other hand, proposes a pruning method specifically designed for Latent Diffusion Models (LDMs) The key innovation of LD-Pruner lies in its utilization of the latent space to guide the pruning process. The method enables a precise assessment of pruning impacts by generating multiple sets of latent vectors—one set for the original Unet and additional sets for each modified Unet where a single operator is altered. The importance of each operator is then quantified using a specialized formula that considers shifts in both the central tendency and variability of the latent vectors. This approach ensures that the pruning process preserves model performance while adapting to the specific characteristics of LDMs.
Lastly, LAPTOPDiff (Zhang et al., 2024a) introduces a layer-pruning technique aimed at automatically compressing the U-Net architecture of diffusion models. The core of this approach is an effective one-shot pruning criterion, distinguished by its favorable additivity property. This property ensures that the one-shot performance of the pruning is superior to other traditional layer pruning methods and manual layer removal techniques. By framing the pruning problem within the context of combinatorial optimization, LAPTOPDiff simplifies the pruning process while achieving significant performance gains. The proposed method stands out for its ability to provide a robust one-shot pruning solution, offering a clear advantage in compressing diffusion models efficiently.
@@ 4 System-Level Efficiency Optimization
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=left, text centered, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=21em,font=,, where level=2text width=15em,font=,, where level=3text width=24em,font=,, [ System, ver [ Optimized Hardware-Software Co-design, fill=blue!10 [  Chen et al. (2023d), SDA (Yang et al., 2023a),  Choi et al. (2024a), leaf, text width=32.7em ] ] [ Parallel Computing, fill=blue!10 [ DistriFusion (Li et al., 2024a), PipeFusion (Wang et al., 2024a),
SwiftDiffusion (Li et al., 2024b),DiffusionPipe (Tian et al., 2024), leaf, text width=32.7em ] ] [ Caching Technique, fill=blue!10 [ NIRVANA (Agarwal et al., 2024), DeepCache (Ma et al., 2024b),
Block Caching (Wimbauer et al., 2024), L2C (Ma et al., 2024a),
FORA (Selvaraju et al., 2024), MD-DiT (Shen et al., ), $\Delta$-DiT (Chen et al., ), leaf, text width=32.7em ] ] ]